# Entanglement Rényi Negativity of Interacting Fermions from Quantum Monte Carlo Simulations

### v2

@@ Abstract
where $S^{\rm th}_{n}=-(\ln\mathrm{Tr}\rho^{n})/(n-1)$ represents the thermodynamic Rényi entropy, which can be computed by considering the subsystem $A_{2}$ as non-existent in $\mathcal{E}_{n}$. The faithful description of mixed-state entanglement necessitates the subtraction of the thermodynamic Rényi entropy $S_{n}^{\rm th}$ from the Rényi negativity $\mathcal{E}_{n}$. In Fig. 2, we display the variations of the negativity ratio and EE with temperature for three distinct lengths, namely $L=6,10,14$. Here, the subsystem $A_{1}$ is selected to be half of the chain, yielding an equal bipartition. As the temperature rises, the EE increases while the negativity ratio asymptotically diminishes to zero for all lengths. This serves as a compelling physical demonstration of the Rényi negativity ratio. In a generic mixed state, both quantum and classical correlations are present, and effective measurement of mixed-state entanglement should exclusively isolate the quantum correlations [28]. In the specific context of finite-temperature Gibbs states, the classical correlation is simply the thermal fluctuations delineated by the thermodynamic entropy $S_{n}^{\rm th}$. Furthermore, at sufficiently low temperatures, the negativity ratio remains constant and establishes a plateau, the length of which is associated with the finite-size gap $1/L$ [62]. As depicted in Fig. 2, it is evident that with an increase in chain length, the plateau becomes narrower. In summary, the monotonic decay of the negativity ratio with rising temperature signifies a crossover from a quantum entangled state to a classical mixed state.
Figure 2: Quantum-classical crossover. The scaled Rényi negativity ratio $R_{2}/L$ and EE $S_{2}/L$ of the half-filled Hubbard chain under a half-chain bipartition vary as functions of temperature. As the temperature rises, the scaled EE for different lengths increases and converges, indicating a dominance of volume law at high temperatures. Meanwhile, the negativity ratio begins to vanish once the temperature reaches a critical value associated with the finite-size gap $1/L$ [62].
Finite-temperature transition in $t$-$V$ model.—To demonstrate the efficacy of Rényi negativity ratio in detecting finite-temperature phase transition, we further consider the half-filled spinless $t$-$V$ model on a square lattice with periodic boundary conditions [83, 84, 82],
@@
where both the hopping and the interaction involve only nearest neighbors. In the presence of a finite coupling parameter $V$, this model exhibits a charge density wave (CDW) ground state and undergoes a phase transition from the CDW phase to a metallic phase at finite temperature, with critical behavior falling within the 2D Ising universality class [84, 85]. In the following, we focus on a specific coupling strength, $V/t=2$, where the critical temperature is estimated to be approximately $T_{c}/t\approx 1.0$, a value lower than the 2D Ising result of $1.120.56|V/t|=$ 1.12 [84].
@@
This model is also a sign-problem-free model [86, 87, 88, 89]. However, for models with larger dimensions or stronger interaction strengths, the direct sampling of Rényi negativity using Eq. (8) becomes inaccurate and challenging to converge, as a result of the occurrence of spikes [90] or the non-Gaussian distribution of Grover determinants $\det g_{x}=\det[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+(I-G_{\bf{s}_{1}}^{T_{2}^{f}})(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}})]$ [91]. We implement an incremental algorithm for rank-$n$ Rényi negativity, analogous to the controllable incremental algorithm for Rényi EE [92, 93, 91, 94], the spirit of which is to measure $(\det g_{x})^{1/N_{\rm inc}}$ instead of $(\det g_{x})$ to circumvent the sampling of an exponentially small quantity with exponentially large variance [95]. It is important to note that there is a sign ambiguity in the $N_{\rm inc}$-th root. In the Supplementary Material [70], we prove that the Grover determinant $\det g_{x}$ is always real and non-negative for two classes of sign-free models, represented by the Hubbard model and the spinless $t$-$V$ model, respectively.
Figure 3: The finite temperature transition in the spinless $t$-$V$ model is detected by the area-law coefficient of the Rényi negativity ratio as a function of temperature. The geometry of the bipartition is illustrated in the inset on the right. A vertical arrow, colored in gray, indicates the position of the shared peak, with half of it aligning with the transition point determined in previous studies [84, 82]. The left inset shows the linear scaling of the area-law coefficient at the critical point with $\ln L$.
@@
As illustrated in the right inset of Fig. 3, we designate the lower left corner with dimensions $(L/2)\times(L/2)$ as subsystem $A_{1}$, resulting in an area-law coefficient of the Rényi negativity ratio of $R_{2}/L$. The main plot of Fig. 3 presents $R_{2}/L$ as a function of temperature for various system sizes, demonstrating a notably distinct finite-size characteristic compared to the intersection of mutual information [96, 97, 98, 82]. Remarkably, unlike the Hubbard model in Fig. 2 or the previous study on the $2+1\mathrm{D}$ transverse field Ising model [47], the Rényi negativity ratio does not exhibit a monotonic decrease with rising temperature. Instead, for varying lattice sizes, a shared local maximum appears at approximately twice the transition temperature, $2T_{c}/t\approx 2.1$. The inclusion of the prefactor $2$ aligns with the rank of the Rényi negativity ratio under consideration, consistent with earlier discussion on the critical behavior of Rényi negativity [51, 53, 47]. The peak of $R_{2}/L$ exhibits a logarithmic divergence with system size $L$ as shown in the left inset of Fig. 3, while the area law is well preserved far away from the critical point  [^4] This beyond-area-law scaling around the finite temperature critical point is also observed for other values of $V$, a different bipartite geometry, and a different lattice. Refer to the SM for additional complementary plots [70].. Based on symmetry considerations, it was argued that the entanglement negativity inherits the singularity of the specific heat at a finite temperature transition [53, 47], and for the 2D Ising transition, the specific heat has a logarithmic divergence. However, in our fermionic scenario, the quantity showing this divergence is $R_{2}/L$ rather than its temperature derivative. The underlying cause of the logarithmic divergence of $R_{2}/L$ at $2T_{c}$ thus remains an unresolved issue.
@@
Conclusions and outlook.—We showed that the partially transposed density matrix for interacting fermions, akin to the reduced density matrix, can be expanded as a weighted sum of Gaussian states representing free fermions, thereby paving the way for the study of mixed-state entanglement in strongly correlated fermionic systems. This main result was employed to implement an algorithm to compute the rank-$n$ Rényi negativity for interacting fermionic systems within the DQMC framework. We presented the first study of the rank-two Rényi negativity for the half-filled Hubbard chain and the spinless $t$-$V$ model on a square lattice. Remarkably, we found that the area law coefficient of the negativity ratio exhibits a logarithmic singular peak at twice the finite-temperature transition point for all lattice sizes under consideration.
@@
There are several potential future research directions to consider. The first direction is to investigate the finite-temperature entanglement of various interacting fermionic models, especially those with transition points that belong to different universality classes, such as the 3D Hubbard model which owns a transition belonging to $\mathrm{O}(3)$ universality class [100]. Further, exploring the entanglement in other types of mixed states, such as tripartite ground states of topological [57] and gapless systems [39], and measurement-induced mixed states [101, 102] presents an intriguing avenue for further research. Next, exploring the finite-size scaling laws of negativity in interacting fermionic systems, especially the beyond-area-law behaviors that reveal universal properties of quantum systems, could also be intriguing. In particular, our finding of the $L\ln L$ scaling of the negativity ratio at the critical point in the spinless $t$-$V$ model may indicate long-range entanglement contribution [47, 65], which warrants further investigation. Moreover, our results are applicable to the continuous-time QMC method, offering an opportunity to study the mixed-state entanglement of realistic correlated materials through combining with dynamical mean-field theory [82, 103, 104]. For example, to fit into the hybridization expansion algorithm, the bath can be first traced out [78, 66] and then the density matrix of the impurity can be obtained using the Green’s functions and density correlation functions [65, 105]. Another interesting example is to integrate the rank-two Rényi negativity into the interaction expansion algorithm by making use the identity ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho X_{2}(\pi))^{2}]$, where the disorder operator, expressed in the interaction picture as $X_{2}(\tau)\propto\prod_{j\in A_{2},\sigma}(n_{i\sigma}(\tau)-\frac{1}{2})$, is nothing just some additional interaction vertices only at $A_{2}$ subregion. Finally, another proposal for realistic materials is to integrate the Rényi negativity into the constrained-path auxiliary-field QMC method [106, 107, 108, 109] which controls the sign problem.
@@ Acknowledgements.
Acknowledgements.—We thank Tarun Grover for helpful discussions and comments on the draft. This work is supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12274289), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and startup funds from SJTU. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.

### v4

@@ II.2 Quantum-classical crossover in Hubbard chain
where $S^{\rm th}_{n}=-(\ln\mathrm{Tr}\rho^{n})/(n-1)$ denotes the thermodynamic Rényi entropy, which equals $\mathcal{E}_{n}$ for either $A_{1}=A$ or $A_{2}=A$. A faithful description of mixed-state entanglement necessitates the exclusion of the thermodynamic Rényi entropy $S_{n}^{\rm th}$. In Fig. 2C, we display the variations of the RNR and EE with temperature for three distinct lengths, namely $L=6,10,14$. Here, the subsystem $A_{1}$ is chosen to be half of the chain, yielding an equal bipartition. As the temperature rises, the EE increases while the RNR asymptotically diminishes to zero for all lengths. This serves as a compelling physical demonstration of the quantity $R_{n}$. In a generic mixed state, both quantum and classical correlations are present, and effective measurement of mixed-state entanglement should exclusively isolate the quantum correlations [15]. In the specific context of finite-temperature Gibbs states, the classical correlation is simply the thermal fluctuations delineated by the thermodynamic entropy $S_{n}^{\rm th}$. Furthermore, at sufficiently low temperatures, the RNR remains constant and establishes a plateau, the length of which is associated with the finite-size gap $1/L$ [44]. As depicted in Fig. 2C, it is evident that with an increase in chain length, the plateau becomes narrower. In summary, the monotonic decay of the RNR with rising temperature signifies a crossover from a quantum entangled state to a classical mixed state.
II.3 Finite temperature transition in $t$-$V$ model.
To demonstrate the efficacy of the RNR in detecting finite-temperature phase transition, we further consider the half-filled spinless $t$-$V$ model on a square lattice with periodic boundary conditions [54, 55, 53],
@@
where both the hopping and the interaction involve only nearest neighbors (see Fig. 3A for illustration). In the presence of a finite coupling parameter $V$, this model exhibits a charge density wave (CDW) ground state and undergoes a phase transition from the CDW phase to a metallic phase at finite temperature (see Fig. 3B), with critical behavior falling within the 2D Ising universality class [55, 56]. In the following, we focus on a specific coupling strength, $V/t=2$, where the critical temperature was estimated to be $T_{c}/t\approx 1.0$ [55].
@@
This model is also a sign-problem-free model [57, 58, 59, 60]. However, for models with larger dimensions or stronger interaction strengths, the direct sampling of RN using Eq. (9) becomes inaccurate, as a result of the occurrence of spikes [61] or the non-Gaussian distribution of Grover determinants $\det g_{x}=\det[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+(I-G_{\bf{s}_{1}}^{T_{2}^{f}})(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}})]$ [62]. We implement an incremental algorithm for the RN, analogous to the controllable incremental algorithm for EE [63, 64, 62], the spirit of which is to measure $(\det g_{x})^{1/N_{\rm inc}}$ instead of $(\det g_{x})$ to circumvent the sampling of an exponentially small quantity with exponentially large variance (see Methods). It is important to note that there is a sign ambiguity in the $N_{\rm inc}$-th root. In the Supplementary Note 5, we prove that the Grover determinant $\det g_{x}$ is always real and non-negative for two classes of sign-free models, represented by the Hubbard model and the spinless $t$-$V$ model, respectively.
Figure 3: Rank-two RNR in the $t$-$V$ model across a finite-temperature transition point. (a) A schematic illustration of the spinless (or spin-polarized) $t$-$V$ model on a square lattice and the chosen one-quarter bipartition geometry. (b) The phase diagram of the $t$-$V$ model, where the phase boundary points at $V=1$ and $V=2$ are determined through mixed-state entanglement studies. One can find the data of $V=1$ case in the Supplementary Fig. 2. The error bars for these two points are estimated based on the difference between the neighboring temperature points that were calculated. (c) The finite-temperature transition in the spinless $t$-$V$ model is detected by the area-law coefficient of the RNR as a function of temperature. A gray vertical arrow indicates the position of the shared peak, with half of its magnitude aligning with the transition point determined in previous studies [55, 53]. The inset shows the linear scaling of the area-law coefficient at the critical point with $\ln L$. The error bars in (c) represent the standard errors from Monte Carlo sampling.
@@
As illustrated in Fig. 3A, we designate the lower left corner with dimensions $(L/2)\times(L/2)$ as subsystem $A_{1}$, resulting in an area-law coefficient of the RNR of $R_{2}/L$. The main plot of Fig. 3C presents $R_{2}/L$ as a function of temperature for various system sizes, demonstrating a notably distinct finite-size characteristic compared to the intersection of mutual information [65, 53]. Remarkably, unlike the Hubbard model in Fig. 2C or the previous study on the $2+1\mathrm{D}$ transverse field Ising model [30], the RNR does not exhibit a monotonic decrease with rising temperature. Instead, for varying lattice sizes, a shared local maximum appears at approximately twice the transition temperature, $2T_{c}/t\approx 2.1$. The inclusion of the prefactor $2$ aligns with the rank of the RNR under consideration, consistent with earlier discussion on the critical behavior within replica approach [33, 35, 30]. In general, the singularity of $R_{n}$ is anticipated to occur at $T=nT_{c}$ with $T_{c}$ being the physical transition temperature. This expectation arises because $R_{n}$, as indicated by the denominator in its definition in Eq. (10), i.e., $\mathrm{Tr}[e^{-n\beta H}]$, effectively corresponds to a Gibbs state with an effective inverse temperature of $n\beta$. Therefore, we demonstrate that it is possible to quantitatively extract the finite-temperature transition points and the phase diagram of the $t$-$V$ model (see Fig. 3B) from mixed-state entanglement studies. Finally, we briefly highlight the finite-size scaling of the rank-two RNR. As shown in the inset of Fig. 3C, the peaks of $R_{2}/L$ exhibit a logarithmic divergence with system size $L$, while the area law is well-preserved in regions far from the critical point. This beyond-area-law scaling around the finite temperature critical point is also observed for other values of $V$ (namely, $V=0$ and $V=1$), a different bipartite geometry, and a different lattice. Refer to the Supplementary Fig. 2 for additional complementary plots.
III Discussion
@@
We showed that the PTDM for interacting fermions, akin to the reduced density matrix, can be expanded as a weighted sum of Gaussian states representing free fermions, thereby paving the way for the study of mixed-state entanglement in strongly correlated fermionic systems. This main result was employed to implement an algorithm to compute the rank-$n$ RN for interacting fermionic systems within the DQMC framework. We studied the rank-two RN for the half-filled Hubbard chain and the spinless $t$-$V$ model on a square lattice. Remarkably, we found that the area law coefficient of the RNR exhibits a logarithmic singular peak at about twice the finite-temperature transition point for all lattice sizes under consideration.
We now discuss the possible physical interpretations of the logarithmic divergence of the rank-two RNR at $2T_{c}$. Based on symmetry considerations, it was argued that the entanglement negativity inherits the singularity of the specific heat at a finite temperature transition [35, 30], and for the 2D Ising transition, the specific heat has a logarithmic divergence in lattice linear size. However, in our fermionic scenario, the quantity showing this divergence is $R_{2}/L$ rather than its temperature derivative. Thus it can not be directly connected to specific heat and the underlying cause of the logarithmic divergence of $R_{2}/L$ at $2T_{c}$ remains an open issue. We note that Refs. [35, 30] concerning bosonic models used the conventional partial transpose, which may partly account for the inconsistency.
@@
There are several potential future research directions to consider. The first direction is to investigate the finite-temperature entanglement of various interacting fermionic models, especially those with transition points that belong to different universality classes, such as the 3D Hubbard model which owns a transition belonging to $\mathrm{O}(3)$ universality class [66]. Further, exploring the entanglement in other types of mixed states, such as tripartite ground states of topological [39] and gapless systems [25], and measurement-induced mixed states [67] presents an intriguing avenue for further research. Next, exploring the finite-size scaling laws of negativity in interacting fermionic systems could also be intriguing. In particular, our finding of the $L\ln L$ scaling of the RNR at the critical point in the spinless $t$-$V$ model may indicate long-range entanglement contribution [30, 68], which warrants further investigation. Moreover, our results are applicable to the continuous-time QMC method, offering an opportunity to study the mixed-state entanglement of realistic correlated materials through combining with dynamical mean-field theory [69, 70, 71, 72, 73]. In the hybridization expansion algorithm, the bath can be firstly traced out [51, 46], allowing the impurity’s reduced density matrix to be derived from the Green’s functions and density correlation functions [68]. Additionally, the rank-two RN can be computed within the interaction expansion algorithm via the identity ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho X_{2}(\pi))^{2}]$. Here, the disorder operator $X_{2}(\tau)\propto\prod_{j\in A_{2},\sigma}(n_{i\sigma}(\tau)-\frac{1}{2})$ in the interaction picture introduces additional interaction vertices exclusively in the $A_{2}$ subregion. Finally, another proposal for realistic materials is to integrate the RN into the constrained-path auxiliary-field QMC method [74, 75, 76, 77] which controls the sign problem.
IV Methods
IV.1 Determinantal Quantum Monte Carlo
We used determinantal qauntum Monte Carlo (QMC) to simulate the two interacting fermionic models, both in zero-temperature regime (i.e., projective QMC) and in finite-temperature regime. Interesting readers may refer to the supplementary materials for all the details including basic formalism and the Hubbard-Stratonovich transformations used in this work. The projective DQMC calculations performed in Fig. 2 used a projection length $\Theta/t=20$, which is long enough to project the trial state to the ground state and ensures desired convergence. We chose the time slice step $\Delta_{\tau}$ to be between $0.02$ and $0.05$, depending on the size of $\Theta$ or $\beta$, and the results do not change if we choose a smaller $\Delta_{\tau}$. The results shown in Fig. 2 of the main text were accelerated by employing the delay update algorithm [78].
IV.2 Incremental algorithm for Rényi negativity
The results presented in Fig. 3 were calculated using the incremental algorithm, which was proposed and implemented for Rényi entanglement entropy [62]. We have developed an analogous version for Rényi negativity. Specifically, we measure the exponentially small observable $e^{-(n-1)\mathcal{E}_{n}}$ by separately calculating its $N_{\rm inc}$ factors, each of which is of order $O(10^{-1})$,
$e^{-(n-1)\mathcal{E}_{n}}$ $=\frac{\sum_{\mathbf{s}}w_{\mathbf{s}_{1}}\cdots w_{\mathbf{s}_{n}}\det g_{x}}{\sum_{\mathbf{s}}w_{\mathbf{s}_{1}}\cdots w_{\mathbf{s}_{n}}}$ (12)
$=\frac{Z_{N_{{\rm inc}}}}{Z_{N_{{\rm inc}}-1}}\cdots\frac{Z_{k+1}}{Z_{k}}\cdots\frac{Z_{1}}{Z_{0}},$
where $w_{\mathbf{s}}$ is the regular DQMC weight for a specific auxiliary field configuration $\mathbf{s}$. Here, we define intermediate partition functions as $Z_{k}=\sum_{{\bf s}_{1}\cdots{\bf s}_{n}}w_{{\bf s}_{1}}\cdots w_{{\bf s}_{n}}\left(\det g_{x}\right)^{\frac{k}{N_{\rm inc}}}$. Each ratio can be interpreted as the average value of the $N_{\rm inc}$-th root of the Grover determinant, $(\det g_{x})^{1/N_{\rm inc}}$, over a replicated system with weight $W_{\mathbf{s}_{1}\cdots\mathbf{s}_{n}}=w_{{\bf s}_{1}}\cdots w_{{\bf s}_{n}}\left(\det g_{x}\right)^{\frac{k}{N_{\rm inc}}}$
$\frac{Z_{k+1}}{Z_{k}}$ $=\frac{\sum_{{\bf s}_{1}\cdots{\bf s}_{n}}w_{{\bf s}_{1}}\cdots w_{{\bf s}_{n}}\left(\det g_{x}\right)^{\frac{k+1}{N_{{\rm inc}}}}}{\sum_{{\bf s}_{1}\cdots{\bf s}_{n}}w_{{\bf s}_{1}}\cdots w_{{\bf s}_{n}}\left(\det g_{x}\right)^{\frac{k}{N_{{\rm inc}}}}}$ (13)
$=\frac{\sum_{{\bf s}_{1}\cdots{\bf s}_{n}}W_{{\bf s}_{1}\cdots{\bf s}_{n}}\left(\det g_{x}\right)^{\frac{1}{N_{{\rm inc}}}}}{\sum_{{\bf s}_{1}\cdots{\bf s}_{n}}W_{{\bf s}_{1}\cdots{\bf s}_{n}}}.$
In our calculations, the number of intermediate processes is $N_{\rm inc}=64$, which is sufficiently large to achieve desirable statistical accuracy for the lattice sizes considered.
V Data and code availability
Data availability
The data that support the findings of this study are provided at https://scidata.sjtu.edu.cn/records/2xm0w-eng25.
Code availability
The code used in this work is available from the corresponding author upon request.
VI Acknowledgements
@@ Acknowledgements.
The authors thank Tarun Grover for helpful discussions and comments on the draft. This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12447103, No. 12274289), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and startup funds from SJTU. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.